OpenStudy (anonymous):

6x^3-14x^2-12x factor trinomial completely

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

take the common, is x

OpenStudy (raden):

so, it can be x(6x^2 - 14x -12)

OpenStudy (raden):

sorry, i think the common should 2x, therefore = 2x(3x^2 - 7x - 6)

OpenStudy (anonymous):

what else is common?

OpenStudy (anonymous):

now factor within ( )

OpenStudy (raden):

now, factor out which in bracket (3x^2 - 7x - 6) = (3x + 2)(x - 3)

OpenStudy (raden):

so, 6x^3-14x^2-12x = 2x(3x + 2)(x - 3)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

2x(3x+2)(x-3)

OpenStudy (anonymous):

2x(3x+2)(x-3)

OpenStudy (raden):

yes, or u can modif again... 2x(3x+2)(x-3) = x(6x+4)(x-3) = x(3x+2)(2x-6) they are is same :)

OpenStudy (anonymous):

all these would be factored completely?

OpenStudy (raden):

actually, yes just choice in ur options there